# A variational discretization method for mean curvature flows by the Onsager principle

### v1

@@ 4.1 Derivation of the dynamic equation by the Onsager principle
For simplicity in presentation, we still consider the two dimensional case in this section. We suppose that the area enclosed by a closed curve $\Gamma(t)$ is $S$. The changing rate of the area with respect to time is calculated by
$\frac{\text{d}S}{\text{d}t}=\int_{\Gamma(t)}\vec{v}\cdot\vec{n}\text{d}s=\int_{\Gamma(t)}v_{n}\text{d}s.$ (33)
We assume that both the energy and the dissipative function are the same as in the section 3.1. The Rayleighian $\mathcal{R}$ is defined in (4). When the area $S$ is preserved, the Onsager principle can be stated as follows. We minimize the Rayleighian $\mathcal{R}$ with respect to $v_{n}$ under the constraint $\frac{\text{d}S}{\text{d}t}=0$, i.e.
$\min_{v_{n}}\ \mathcal{R},$ (34)
@@
$\mathcal{L}(v_{n},\lambda)=\mathcal{R}+\lambda\frac{\text{d}S}{\text{d}t}=\int_{\Gamma(t)}(\frac{1}{2}v_{n}^{2}+\kappa v_{n}+\lambda v_{n})\text{d}s.$ (35)
@@
$\left\{\begin{array}[]{rcl}\frac{\partial\mathcal{L}}{\partial v_{n}}=0\\ \frac{\partial\mathcal{L}}{\partial\lambda}=0.\end{array}\right.$ (36)
@@
$\left\{\begin{array}[]{rcl}v_{n}=-(\kappa+\lambda)\\ \int_{\Gamma(t)}v_{n}\text{d}s=0.\end{array}\right.$ (37)
The equation can be simplified as follows. We substitute the first equation of (37) into the second equation. This leads to
$\int_{\Gamma(t)}-(\kappa+\lambda)\text{d}s=0,$ (38)
@@
$\lambda|\Gamma(t)|=-\int_{\Gamma(t)}\kappa\text{d}s.$ (39)
@@
$\lambda=-\frac{2\pi}{|\Gamma(t)|}.$ (41)
Therefore, we obtain the normal velocity of the curve $\Gamma(t)$
$v_{n}=-(\kappa-\frac{2\pi}{|\Gamma(t)|}).$ (42)
By substituting the equation (42) into (10), we have
$\dot{\mathcal{E}}$ $=\int_{\Gamma(t)}\kappa v_{n}\text{d}s=-\int_{\Gamma(t)}\kappa(\kappa-\frac{2\pi}{|\Gamma(t)|})\text{d}s=-\int_{\Gamma(t)}(\kappa-\frac{2\pi}{|\Gamma(t)|})^{2}\text{d}s=-2\Phi(v_{n})\leq 0.$ (43)
Here we have used the Gauss-Bonnet formula (40) so that
$\int_{\Gamma(t)}\frac{2\pi}{|\Gamma(t)|}(\kappa-\frac{2\pi}{|\Gamma(t)|})\text{d}s=\frac{2\pi}{|\Gamma(t)|}\int_{\Gamma(t)}(\kappa-\frac{2\pi}{|\Gamma(t)|})\text{d}s=0.$
@@
Similarly to the case of standard mean curvature flow, the equation (42) is equivalent to the following equation
$\vec{v}=-(\kappa-\frac{2\pi}{|\Gamma(t)|})\vec{n}.$ (44)
@@ 4.2 Discretization of the dynamic equation by the Onsager principle
In this subsection, we will numerically solve the dynamic equation (37) of $\Gamma(t)$ for the volume preserving mean curvature flow. The discretization of the curve is the same as that in Section 3.2 (Figure 1).
@@
$\frac{\text{d}S_{h}}{\text{d}t}=\int_{\Gamma_{h}(t)}\vec{v}_{h}(t)\cdot\vec{n}_{i}\text{d}s=\sum_{i=1}^{n}\int_{\Gamma_{i}(t)}\vec{v}_{h}(t)\cdot\vec{n}_{i}\text{d}s.$ (45)
@@
$\mathcal{L}_{h}(\dot{\vec{x}}_{i},\lambda)=\mathcal{R}_{h}+\lambda\frac{\text{d}S_{h}}{\text{d}t}.$ (47)
@@
$\left\{\begin{array}[]{rcl}\frac{\partial\mathcal{L}_{h}}{\partial\dot{\vec{x}}_{i}}=0\\ \frac{\partial\mathcal{L}_{h}}{\partial\lambda}=0.\end{array}\right.$ (48)
@@
$\left\{\begin{array}[]{rcl}\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}+\lambda\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}=0\\ \frac{\text{d}S_{h}}{\text{d}t}=0.\end{array}\right.$ (49)
@@
Denote by $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T},\lambda)^{T}$ and notice that $\Phi_{h}$ is a quadratic function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. The system (49) can be rewritten as
@@
where the coefficient matrix $\hat{A}$ is a symmetric positive definite matrix and $\hat{G}=(\vec{g}_{1}^{T},\vec{g}_{2}^{T},...,\vec{g}_{n}^{T},g_{n+1})^{T}$. The explicit formulas for $\hat{A}$ and $\hat{G}$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T},\lambda)^{T}$ is the solution of the equation (50), then we have
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=-2\Phi_{h}\leq 0,$ (51)
@@
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}$ $=\sum_{i=1}^{n}\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}\cdot\dot{\vec{x}}_{i}=-\sum_{i=1}^{n}[\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\lambda\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}]\cdot\dot{\vec{x}}_{i}$ (52)
$=-\sum_{i=1}^{n}\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}\cdot\dot{\vec{x}}_{i}-\lambda\sum_{i=1}^{n}\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}\cdot\dot{\vec{x}}_{i}=-\dot{X}_{0}^{T}A_{0}\dot{X}_{0}-\lambda\frac{\text{d}S_{h}}{\text{d}t}$
$=-\dot{X}_{0}^{T}A_{0}\dot{X}_{0}=-2\Phi_{h}.$
Here we have used the first equation of (49) and $\frac{\text{d}S_{h}}{\text{d}t}=0$. It is also easy to know that the matrix $A_{0}$ is symmetric positive definite, so the theorem is proved.
@@ 5 Wetting problems
In this section, we further apply the method to wetting problems which can be formulated as a volume conserving mean curvature flow for curves with contact with the substrates.
@@ 5.1 Derivation of the dynamic equation by the Onsager principle
[…]
@@
[…]
$\gamma_{LV}\cos\theta_{Y}=\gamma_{SV}-\gamma_{SL}.$ (54)
Figure 2: Schematic diagram of a liquid-vapor-solid system
@@
~$\tilde{\mathcal{E}}=\gamma_{LV}|\Gamma_{LV}|+\gamma_{SL}|\Gamma_{SL}|+\gamma_{SV}|\Gamma_{SV}|=\gamma_{LV}|\Gamma_{LV}|-\gamma_{LV}\cos\theta_{Y}|\Gamma_{SL}|+C,$ (55)
@@
Consider a general two dimensional non-closed curve $\tilde{\Gamma}(t)$, where the left and right endpoints lie on the same horizontal line, as shown in Figure 3. Let $\vec{x}(s,t)$ denote the points on the curve $\tilde{\Gamma}(t)$, where $s$ is the arc length parameter. We define the counterclockwise direction as the positive direction of the curve $\tilde{\Gamma}(t)$, and denote its length as $L(t)$. The right endpoint is represented by $\vec{x}(0,t)$ and the left endpoint by $\vec{x}(L,t)$. The left and right contact angles between the curve $\tilde{\Gamma}(t)$ and the horizontal line are denoted by $\theta_{L}$ and $\theta_{0}$, respectively. The unit tangent vector is denoted by $\vec{\tau}$ and the outward unit normal vector by $\vec{n}$. Under the constraint that the area enclosed by $\tilde{\Gamma}(t)$ and the horizontal line remains constant, we aim to derive the evolution equation of the non-closed curve $\tilde{\Gamma}(t)$ under mean curvature flow.
Figure 3: A non-closed curve whose two endpoints always lie on the same horizontal line
@@
~$\tilde{\mathcal{E}}=\gamma[L(t)-|\vec{x}(0,t)-\vec{x}(L,t)|\cos\theta_{Y}]+C,$ (56)
[…]
$\frac{\text{d}L(t)}{\text{d}t}=\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v(0)\cos\theta_{0}-v(L)\cos\theta_{L}.$ (57)
@@
$\dot{\tilde{\mathcal{E}}}$ $=\gamma[\frac{\text{d}L(t)}{\text{d}t}-\frac{\text{d}}{\text{d}t}|\vec{x}(0,t)-\vec{x}(L,t)|\cos\theta_{Y}]$ (58)
$=\gamma[\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v(0)\cos\theta_{0}-v(L)\cos\theta_{L}-\cos\theta_{Y}(v(0)-v(L))]$
$=\gamma[\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v(0)(\cos\theta_{0}-\cos\theta_{Y})-v(L)(\cos\theta_{L}-\cos\theta_{Y})].$
@@
$\tilde{\Phi}=\frac{\xi_{0}}{2}\int_{\tilde{\Gamma}(t)}v_{n}^{2}\text{d}s+\frac{\xi_{1}}{2}v(0)^{2}+\frac{\xi_{1}}{2}v(L)^{2},$ (59)
where $\xi_{0},\xi_{1}$ are positive friction coefficients. Notice we introduce additional terms $\frac{\xi_{1}}{2}v(0)^{2}$ and $\frac{\xi_{1}}{2}v(L)^{2}$ to the dissipation function $\tilde{\Phi}$, which are dissipations due to the contact line frictions. As a result, it characterize the relaxation process of the contact angles from their initial values $\theta_{L}$ and $\theta_{0}$ to the equilibrium Young’s angle $\theta_{Y}$. Then the corresponding Rayleighian is given by
~$\tilde{\mathcal{R}}=\tilde{\Phi}+\dot{\tilde{\mathcal{E}}}.$ (60)
[…]
$\frac{\text{d}\tilde{S}}{\text{d}t}=\int_{\tilde{\Gamma}(t)}v_{n}\text{d}s.$ (61)
By using the Onsager principle, we minimize the Rayleighian $\tilde{\mathcal{R}}$ with respect to $v_{n},v(0),v(L)$ under the area conservation condition, i.e.
$\min_{v_{n},v(0),v(L)}\ \tilde{\mathcal{R}},$ (62)
@@
~$\tilde{\mathcal{L}}(v_{n},v(0),v(L),\lambda)=\tilde{\mathcal{R}}+\lambda\frac{\text{d}\tilde{S}}{\text{d}t}$ $=\int_{\tilde{\Gamma}(t)}(\frac{\xi_{0}}{2}v_{n}^{2}+\gamma\kappa v_{n}+\lambda v_{n})\text{d}s+\frac{\xi_{1}}{2}v(0)^{2}+\frac{\xi_{1}}{2}v(L)^{2}$ (63)
$\quad+\gamma[v(0)(\cos\theta_{0}-\cos\theta_{Y})-v(L)(\cos\theta_{L}-\cos\theta_{Y})].$
@@
{~$\left\{\begin{array}[]{rcl}\frac{\partial\tilde{\mathcal{L}}}{\partial v_{n}}=0\\ \frac{\partial\tilde{\mathcal{L}}}{\partial v(0)}=0\\ \frac{\partial\tilde{\mathcal{L}}}{\partial v(L)}=0\\ \frac{\partial\tilde{\mathcal{L}}}{\partial\lambda}=0.\end{array}\right.$ (64)
@@
$\left\{\begin{array}[]{rcl}&\xi_{0}v_{n}=-(\gamma\kappa+\lambda)\\ &\xi_{1}v(0)=-\gamma(\cos\theta_{0}-\cos\theta_{Y})\\ &\xi_{1}v(L)=\gamma(\cos\theta_{L}-\cos\theta_{Y})\\ &\int_{\tilde{\Gamma}(t)}v_{n}\text{d}s=0.\end{array}\right.$ (65)
Substitute the first equation of (65) into the last one. We get
$\int_{\tilde{\Gamma}(t)}-(\gamma\kappa+\lambda)\text{d}s=0.$ (66)
@@
$\int_{\tilde{\Gamma}(t)}\kappa\text{d}s+0\cdot|\vec{x}(0,t)-\vec{x}(L,t)|+\pi-\theta_{0}+\pi-\theta_{L}=2\pi.$ (68)
@@
Substituting the equation (69) into (67), we have
$\lambda=-\frac{\gamma(\theta_{0}+\theta_{L})}{L(t)}.$ (70)
Thus, we obtain the normal velocity of the curve $\tilde{\Gamma}(t)$
$\xi_{0}v_{n}=-\gamma(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)}).$ (71)
Substituting the equation (65) and the equation (71) into (58), we get
$\dot{\tilde{\mathcal{E}}}$ $=\gamma[\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v(0)(\cos\theta_{0}-\cos\theta_{Y})-v(L)(\cos\theta_{L}-\cos\theta_{Y})]$ (72)
$=\gamma[\int_{\tilde{\Gamma}(t)}-\xi_{0}^{-1}\gamma\kappa(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)})\text{d}s-\xi_{1}^{-1}\gamma(\cos\theta_{0}-\cos\theta_{Y})^{2}-\xi_{1}^{-1}\gamma(\cos\theta_{L}-\cos\theta_{Y})^{2}]$
$=-\gamma^{2}[\int_{\tilde{\Gamma}(t)}\xi_{0}^{-1}(\kappa^{2}-\kappa\frac{\theta_{0}+\theta_{L}}{L(t)})\text{d}s+\xi_{1}^{-1}(\cos\theta_{0}-\cos\theta_{Y})^{2}+\xi_{1}^{-1}(\cos\theta_{L}-\cos\theta_{Y})^{2}].$
@@
$\int_{\tilde{\Gamma}(t)}\kappa(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)})\text{d}s=\int_{\tilde{\Gamma}(t)}(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)})^{2}\text{d}s.$

### v3

@@ 4.1 Derivation of the dynamic equation by the Onsager principle
We suppose that the area enclosed by a closed curve $\Gamma(t)$ is $S$. The changing rate of the area with respect to time is calculated by
$\frac{\text{d}S}{\text{d}t}=\int_{\Gamma(t)}\vec{v}\cdot\vec{n}\text{d}s=\int_{\Gamma(t)}v_{n}\text{d}s.$
We assume that both the energy and the dissipative function are the same as in the section 3.1. The Rayleighian $\mathcal{R}$ is defined in (2). When the area $S$ is preserved, the Onsager principle can be stated as follows. We minimize the Rayleighian $\mathcal{R}$ with respect to $\vec{v}$ under the constraint $\frac{\text{d}S}{\text{d}t}=0$, i.e.
$\min_{\vec{v}}\ \mathcal{R},$ (18)
@@
$\mathcal{L}(\vec{v},\lambda)=\mathcal{R}+\lambda\frac{\text{d}S}{\text{d}t}=\int_{\Gamma(t)}\left(\frac{1}{2}|\vec{v}|^{2}+\kappa v_{n}+\lambda v_{n}\right)\text{d}s.$
@@
$\left\{\begin{array}[]{rcl}\frac{\delta\mathcal{L}}{\delta\vec{v}}=0,\vspace{1ex}\\ \frac{\partial\mathcal{L}}{\partial\lambda}=0.\end{array}\right.$
@@
$\vec{v}=-(\kappa+\lambda)\vec{n},$ (19a)
$\int_{\Gamma(t)}v_{n}\text{d}s=0.$ (19b)
The equation can be simplified as follows. We substitute Eq. (19a) into (19b). This leads to
$\int_{\Gamma(t)}-(\kappa+\lambda)\text{d}s=0,$
@@
$\lambda|\Gamma(t)|=-\int_{\Gamma(t)}\kappa\text{d}s.$
@@
$\lambda=-\frac{2\pi}{|\Gamma(t)|}.$
Therefore, we obtain the equation of volume preserving mean curvature flow as follows:
$\vec{v}=-\left(\kappa-\frac{2\pi}{|\Gamma(t)|}\right)\vec{n}.$ (21)
By substituting Eq. (21) into (8), we have
˙$\dot{\mathcal{E}}=\int_{\Gamma(t)}\kappa v_{n}\text{d}s=-\int_{\Gamma(t)}\kappa\left(\kappa-\frac{2\pi}{|\Gamma(t)|}\right)\text{d}s=-\int_{\Gamma(t)}\left(\kappa-\frac{2\pi}{|\Gamma(t)|}\right)^{2}\text{d}s=-2\Phi\leq 0.$
Here we have used Eq. (20) so that
$\int_{\Gamma(t)}\frac{2\pi}{|\Gamma(t)|}\left(\kappa-\frac{2\pi}{|\Gamma(t)|}\right)\text{d}s=\frac{2\pi}{|\Gamma(t)|}\left(\int_{\Gamma(t)}\kappa\text{d}s-\frac{2\pi}{|\Gamma(t)|}\int_{\Gamma(t)}\text{d}s\right)=0.$
@@ 4.2 Discretization of the dynamic equation by the Onsager principle
In this subsection, we will numerically solve the dynamic equation (19b) of $\Gamma(t)$ for the volume preserving mean curvature flow. The discretization of the curve is the same as that in Section 3.2 (Fig. 1).
@@
$\frac{\text{d}S_{h}}{\text{d}t}=\int_{\Gamma_{h}(t)}v_{h,n}\text{d}s=\sum_{i=1}^{n}\int_{\Gamma_{i}(t)}\vec{v}_{h}\cdot\vec{n}_{i}\text{d}s.$
@@
$\mathcal{L}_{h}(\dot{\vec{x}}_{i},\lambda)=\mathcal{R}_{h}+\lambda\frac{\text{d}S_{h}}{\text{d}t}.$
@@
$\left\{\begin{array}[]{rcl}\frac{\partial\mathcal{L}_{h}}{\partial\dot{\vec{x}}_{i}}=0,\vspace{1ex}\\ \frac{\partial\mathcal{L}_{h}}{\partial\lambda}=0.\end{array}\right.$
@@
$\left\{\begin{array}[]{ll}\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\lambda\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}=-\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}},\vspace{1ex}\\ \frac{\text{d}S_{h}}{\text{d}t}=0.\end{array}\right.$ (23)
@@
Denote by $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top},\lambda)^{\top}$ and notice that $\Phi_{h}$ is a quadratic function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. The system (23) can be rewritten as
@@
where the coefficient matrix $\hat{A}$ is a symmetric matrix and $\hat{G}=(\vec{g}_{1}^{\top},\vec{g}_{2}^{\top},...,\vec{g}_{n}^{\top},g_{n+1})^{\top}$. The explicit formulas for $\hat{A}$ and $\hat{G}$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top},\lambda)^{\top}$ is the solution of Eq. (24), then we have
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}=-2\Phi_{h}\leq 0,$
@@
$\frac{\text{d}\mathcal{E}_{h}}{\text{d}t}$ $=\sum_{i=1}^{n}\frac{\partial\mathcal{E}_{h}}{\partial\vec{x}_{i}}\cdot\dot{\vec{x}}_{i}=-\sum_{i=1}^{n}\left[\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}+\lambda\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}\right]\cdot\dot{\vec{x}}_{i}$
$=-\sum_{i=1}^{n}\frac{\partial\Phi_{h}}{\partial\dot{\vec{x}}_{i}}\cdot\dot{\vec{x}}_{i}-\lambda\sum_{i=1}^{n}\frac{\partial(S_{h})_{t}}{\partial\dot{\vec{x}}_{i}}\cdot\dot{\vec{x}}_{i}=-\dot{X}_{0}^{\top}A_{0}\dot{X}_{0}-\lambda\frac{\text{d}S_{h}}{\text{d}t}$
$=-\dot{X}_{0}^{\top}A_{0}\dot{X}_{0}=-2\Phi_{h}.$
Here we have used Eq. (23). It is also easy to know that the matrix $A_{0}$ is symmetric positive definite, so the theorem is proved.
In simulations, we also use the modified energy function $\mathcal{E}_{h}^{\delta}$ to avoid the mesh degeneracy, as in section 3.3, and the corresponding Euler-Lagrange equation is
$\hat{A}\dot{X}=\hat{G}^{\delta}.$ (25)
@@ 5 Wetting problems
In this section, we further apply the method to wetting problems which can be formulated as a volume preserving mean curvature flow for curves in contact with the substrates.
@@ 5.1 Derivation of the dynamic equation by the Onsager principle
We consider a two dimensional wetting problem, as illustrated in Fig. 2, where the total energy of a liquid-vapor-solid system is composed of three components (c.f. [50]), i.e.
@@
where $\gamma_{LV}$, $\gamma_{SL}$ and $\gamma_{SV}$ represent the energy densities of the liquid-vapor interface $\Gamma_{LV}$, the solid-liquid interface $\Gamma_{SL}$ and the solid-vapor interface $\Gamma_{SV}$ respectively. Assuming that the solid boundary $\Gamma_{S}=\Gamma_{SL}\bigcup\Gamma_{SV}$ is homogeneous, then both $\gamma_{SL}$ and $\gamma_{SV}$ are constants. In equilibrium the angle between the liquid-vapor interface and the solid interface is determined by Young’s equation [51]:
$\gamma_{LV}\cos\theta_{Y}=\gamma_{SV}-\gamma_{SL}.$
@@
$\tilde{\mathcal{E}}$ $=\gamma_{LV}|\Gamma_{LV}|-\frac{\gamma_{LV}\cos\theta_{Y}}{2}|\Gamma_{SL}|+\frac{\gamma_{LV}\cos\theta_{Y}}{2}(|\Gamma_{S}|-|\Gamma_{SL}|)+\frac{\gamma_{SL}+\gamma_{SV}}{2}|\Gamma_{S}|$ (27)
$=\gamma_{LV}|\Gamma_{LV}|-\gamma_{LV}\cos\theta_{Y}|\Gamma_{SL}|+C,$
@@
Figure 2: Schematic diagram of a liquid-vapor-solid system.
Consider a general non-closed curve $\tilde{\Gamma}(t)$, where the left and right endpoints lie on the same horizontal line, as shown in Fig. 3. Let $\vec{x}(s,t)$ denote the points on the curve $\tilde{\Gamma}(t)$, where $s$ is the arc length parameter. We define the counterclockwise direction as the positive direction of the curve $\tilde{\Gamma}(t)$, and denote its length as $L(t)$. The right endpoint is represented by $\vec{x}(0,t)$ and the left endpoint by $\vec{x}(L,t)$. The left and right contact angles between the curve $\tilde{\Gamma}(t)$ and the horizontal line are denoted by $\theta_{L}$ and $\theta_{0}$, respectively. The unit tangent vector is denoted by $\vec{\tau}$ and the outward unit normal vector by $\vec{n}$. Under the constraint that the area enclosed by $\tilde{\Gamma}(t)$ and the horizontal line remains constant, we aim to derive the evolution equation of the non-closed curve $\tilde{\Gamma}(t)$ under volume preserving mean curvature flow.
Figure 3: A non-closed curve whose two endpoints always lie on the same horizontal line.
@@
~$\tilde{\mathcal{E}}=\gamma[L(t)-|\vec{x}(0,t)-\vec{x}(L,t)|\cos\theta_{Y}]+C,$
where $\gamma=\gamma_{LV}$ represents the surface energy density and $\theta_{Y}$ denotes the Young’s angle. Since the left and right endpoints of $\tilde{\Gamma}(t)$ lie on the bottom line, they possess solely horizontal velocities. Taking the horizontal right direction as the positive direction, the velocities at the left and right endpoints are denoted by $v_{L}(t)$ and $v_{0}(t)$, respectively. The changing rate of the length $L(t)$ of $\tilde{\Gamma}(t)$ with respect to time is given by (c.f. Lemma 5.1 in [36])
$\frac{\text{d}L(t)}{\text{d}t}=\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v_{0}\cos\theta_{0}-v_{L}\cos\theta_{L}.$
@@
$\dot{\tilde{\mathcal{E}}}$ $=\gamma\left[\frac{\text{d}L(t)}{\text{d}t}-\frac{\text{d}}{\text{d}t}|\vec{x}(0,t)-\vec{x}(L,t)|\cos\theta_{Y}\right]$ (28)
$=\gamma\left[\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v_{0}\cos\theta_{0}-v_{L}\cos\theta_{L}-\cos\theta_{Y}(v_{0}-v_{L})\right]$
$=\gamma\left[\int_{\tilde{\Gamma}(t)}\kappa v_{n}\text{d}s+v_{0}(\cos\theta_{0}-\cos\theta_{Y})-v_{L}(\cos\theta_{L}-\cos\theta_{Y})\right].$
@@
$\tilde{\Phi}=\frac{\xi_{0}}{2}\int_{\tilde{\Gamma}(t)}|\vec{v}|^{2}\text{d}s+\frac{\xi_{1}}{2}v_{0}^{2}+\frac{\xi_{1}}{2}v_{L}^{2},$ (29)
where $\xi_{0},\xi_{1}$ are positive friction coefficients. Notice we introduce additional terms $\frac{\xi_{1}}{2}v_{0}^{2}$ and $\frac{\xi_{1}}{2}v_{L}^{2}$ to the dissipation function $\tilde{\Phi}$, which are dissipations due to the contact line frictions. As a result, it characterize the relaxation process of the contact angles from their initial values $\theta_{L}$ and $\theta_{0}$ to the equilibrium Young’s angle $\theta_{Y}$. Then the corresponding Rayleighian is given by
~$\tilde{\mathcal{R}}=\tilde{\Phi}+\dot{\tilde{\mathcal{E}}}.$
In wetting problems, the volume of a liquid is preserved. The time derivative of the area $\tilde{S}$ enclosed by the curve $\tilde{\Gamma}(t)$ and the bottom line is given by (c.f. Lemma5.1 in [36])
$\frac{\text{d}\tilde{S}}{\text{d}t}=\int_{\tilde{\Gamma}(t)}v_{n}\text{d}s.$
By using the Onsager principle, we minimize the Rayleighian $\tilde{\mathcal{R}}$ with respect to $\vec{v}$, $v_{0}$, $v_{L}$ under the area conservation condition, i.e.
$\min_{\vec{v},v_{0},v_{L}}\ \tilde{\mathcal{R}},$
@@
~$\tilde{\mathcal{L}}(\vec{v},v_{0},v_{L},\lambda)=\tilde{\mathcal{R}}+\lambda\frac{\text{d}\tilde{S}}{\text{d}t}$ $=\int_{\tilde{\Gamma}(t)}\left(\frac{\xi_{0}}{2}|\vec{v}|^{2}+\gamma\kappa v_{n}+\lambda v_{n}\right)\text{d}s+\frac{\xi_{1}}{2}v_{0}^{2}+\frac{\xi_{1}}{2}v_{L}^{2}$
$\quad+\gamma[v_{0}(\cos\theta_{0}-\cos\theta_{Y})-v_{L}(\cos\theta_{L}-\cos\theta_{Y})].$
@@
$\left\{\begin{array}[]{rcl}\frac{\delta\tilde{\mathcal{L}}}{\delta\vec{v}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}}{\partial v_{0}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}}{\partial v_{L}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}}{\partial\lambda}=0.\end{array}\right.$
@@
$\xi_{0}\vec{v}=-(\gamma\kappa+\lambda)\vec{n},$ (30a)
$\xi_{1}v_{0}=-\gamma(\cos\theta_{0}-\cos\theta_{Y}),$ (30b)
$\xi_{1}v_{L}=\gamma(\cos\theta_{L}-\cos\theta_{Y}),$ (30c)
$\int_{\tilde{\Gamma}(t)}v_{n}\text{d}s=0.$ (30d)
Substitute Eq. (30a) into (30d). We get
$\int_{\tilde{\Gamma}(t)}-(\gamma\kappa+\lambda)\text{d}s=0.$
@@
$\int_{\tilde{\Gamma}(t)}\kappa\text{d}s+0\cdot|\vec{x}(0,t)-\vec{x}(L,t)|+\pi-\theta_{0}+\pi-\theta_{L}=2\pi.$
@@
Substituting Eq. (32) into (31), we have
$\lambda=-\frac{\gamma(\theta_{0}+\theta_{L})}{L(t)}.$
Therefore, the dynamic equation for a wetting problem can be described by
$\left\{\begin{array}[]{rcl}&\vec{v}=-\xi_{0}^{-1}\gamma\left(\kappa-\frac{\theta_{0}+\theta_{L}}{|\tilde{\Gamma}(t)|}\right)\vec{n},\vspace{1ex}\\ &v_{0}=-\xi_{1}^{-1}\gamma(\cos\theta_{0}-\cos\theta_{Y}),\vspace{1ex}\\ &v_{L}=\xi_{1}^{-1}\gamma(\cos\theta_{L}-\cos\theta_{Y}),\end{array}\right.$ (33)
where $\vec{v}$ denotes the velocity of the curve, $v_{0}$ and $v_{L}$ are respectively the horizontal velocity of the two endpoints, $\theta_{L}$ and $\theta_{0}$ denote the left and right contact angles of the curve with the bottom line, and $|\tilde{\Gamma}(t)|$ denotes the length of the curve. Substituting Eq. (33) into (28), we get
$\dot{\tilde{\mathcal{E}}}$ $=-\gamma^{2}\left[\int_{\tilde{\Gamma}(t)}\xi_{0}^{-1}\kappa\left(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)}\right)\text{d}s+\xi_{1}^{-1}(\cos\theta_{0}-\cos\theta_{Y})^{2}+\xi_{1}^{-1}(\cos\theta_{L}-\cos\theta_{Y})^{2}\right].$ (34)
@@
$\int_{\tilde{\Gamma}(t)}\kappa\left(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)}\right)\text{d}s=\int_{\tilde{\Gamma}(t)}\left(\kappa-\frac{\theta_{0}+\theta_{L}}{L(t)}\right)^{2}\text{d}s.$